Schedulers

Immediate  RR Round Robin
EF Earliest First
LL Lightest Loaded
Batch  MM Min-Min
MX Max-Min
DL Dynamic Level
RC Relative Cost

Evolutionary PN This paper

ZO Genetic Algorithm
SA Simlulated Annealing